# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ IV.3 Generalization to monopole pairing orders with higher monopole charges
We further demonstrate the probe of global angular momentum for monopole superconductors in different topological sectors characterized by higher monopole charges, $|q_{p}|>1$. Although the topological sector is different, the same symmetry principles to extract $l_{z,\mathrm{gob}}$ outlined in Sec. III.2 still hold. We begin with the general model of the monopole superconductor in Eq. (13) but now consider a multi-Weyl semimetal with inter-Fermi surface pairing between two parity-related Fermi surfaces. For three-dimensional Weyl semimetals on a lattice with $n$-fold rotational symmetry ($n=2,3,4,6$), there can exist single, double, or triple Weyl nodes [38, 39]. The lattice point group symmetry sets an upper bound to the chirality of the Weyl node, with $|\nu|\leq[n/2]$, and similarly for the pair monopole charge, $|q_{p}|\leq n/2$.
@@
in which $\sigma_{\pm}=\left(\sigma_{x}\pm i\sigma_{y}\right)/2$, and is related to the other Fermi surface by parity, $\mathcal{H}_{\mathrm{kin},1}^{(\nu=-3)}(\mathbf{k})=\sigma_{z}\mathcal{H}_{\mathrm{kin},2}^{(\nu=+3)}(-\mathbf{k})\sigma_{z}$. The helical eigenstate at $\mathrm{FS}_{1}$ is $\chi_{1,+}(\tilde{\mathbf{k}})=(\cos({\theta_{h_{1}}}/{2})e^{-i(\lambda-1)\varphi_{h_{1}}/2},\sin({\theta_{h_{1}}}/{2})e^{-i(\lambda+1)\varphi_{h_{1}}/2})$, and at $\mathrm{FS}_{2}$, it is $\chi_{2,+}(\tilde{\mathbf{k}})=\sigma_{z}\chi_{1,+}(-\tilde{\mathbf{k}})$. Here, consistent with previous notation, $\lambda=\pm 1$ denotes the choice of gauge, and $\theta_{h_{1}}$ and $\varphi_{h_{1}}$ are the polar and azimuthal angles defined in $\mathbf{h}_{1}^{(\nu=-3)}$ spherical coordinates by $\mathbf{h}_{1}^{(\nu=-3)}\equiv|h_{1}^{(\nu=-3)}|(\sin\theta_{h_{1}}\cos\varphi_{h_{1}},\sin\theta_{h_{1}}\sin\varphi_{h_{1}},\cos\theta_{h_{1}})$. They are related to momentum space polar and azimuthal angles by $\sin\theta_{h_{1}}=\sin^{3}\theta_{\tilde{k}}/{(\cos^{2}\theta_{\tilde{k}}+\sin^{6}\theta_{\tilde{k}})^{1/2}}$ and $\varphi_{h_{1}}=-3\varphi_{\tilde{k}}$. We study the case of inter-Fermi surface $s$-wave pairing in Eq. (13), in which $\Delta_{\mathrm{inter}}(\mathbf{k})=i\Delta_{0}\sigma_{y}$, with $\Delta_{0}$ being the pairing amplitude.
In the weak-coupling regime, we obtain the low energy band-projected pairing order from Eq. (16) which takes the following form,
@@
We now examine the Josephson coupling between the above $q_{p}=3$ monopole superconductor and a spin singlet superconductor, following the methods introduced in Sec. III.1. For comparison, we consider the case in which there is bare spin-independent tunneling at the junction barrier. The form factor for this system evaluates to
@@
in which $\bm{\Omega}_{h_{1}}=(\theta_{h_{1}},\varphi_{h_{1}})$ is the solid angle in $\mathrm{h}_{1}^{(\nu=3)}$-space and $\Delta_{\mathrm{MSC}}^{(q_{p}=3}(\tilde{\mathbf{k}})$ is the effective pairing order of the $q_{p}=3$ monopole superconductor. Similar to the case of the $q_{p}=-1$ monopole superconductor, whose form factor is given in Eq. (27), the half-integer monopole harmonics arise from the effective spin-dependent tunneling arising from the spin-orbit coupling in the triple Weyl semimetal. The above form factor is non-singular and can be used to determine the first order Josephson coupling of this system.
@@
in which $\Delta_{L,0}$ and $\Delta_{R,0}$ are the pairing amplitudes of the $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ monopole superconductor respectively. $\mathfrak{F}_{s-\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ itself is invariant under rotation $R_{z}$ and parity, as both the $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ monopole superconductor both form irreps of $l_{z}=0$. As discussed in Sec. IV.1.1, this leads to nonvanishing first order Josephson current. Hence, following the junction design principles introduced in Sec. III.2, we may similarly design probes of monopole pairing order in different topological sectors.
@@ V Josephson junctions to probe $l_{z,\mathrm{loc}}$ of the monopole superconducting order
We now explore how to design Josephson junctions to probe local angular momentum $l_{z,\mathrm{loc}}$ of monopole superconducting order, following the second junction design introduced in Sec. III.2. We consider two identical monopole superconductors forming a Josephson junction oriented along, for example, the $x$ direction, which is perpendicular to the high-symmetry rotational axis along $z$. As discussed in Sec. III.2, the total Josephson current can be regarded as a superposition of contributions from a family of Josephson junctions between two 2D chiral superconductors, corresponding to different values of conserved momentum $k_{z}$. It has been studied extensively in, for example, Refs. 23, 29 that a planar junction between two chiral superconductors can host $4\pi$-periodic Josephson current arising from the tunneling of chiral boundary modes. Therefore, by studying the Josephson energy phase relation in our system, we can obtain the information about local symmetry of monopole pairing order. In the continuum model of a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, its pairing order, described by Eq. (18), exhibits local chiral $p$-wave pairing with $l_{z,\mathrm{loc}}=\mp 1$ for a given value of $\tilde{k}_{z}$,
@@
Figure 5: (a) Single Josephson junction along the $x$ direction between two identical monopole superconductors with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ and the same crystal orientation. (b) Same Josephson junction as in Fig. (a), only now that the superconductor at the left (yellow) is rotated by an angle $\theta$ about the $z$ axis pointing out of plane. (c) Bulk Fermi surface $\mathrm{FS}_{1}$ of the doped Weyl semimetal on a cubic lattice. Representative momenta $\tilde{\mathbf{k}}_{i=1,\cdots,5}$ are shown, with $\tilde{k}_{y}=0$ and $\tilde{k}_{z}=-0.2,-0.1,0.0,0.1,0.15$, respectively. As a reference, we plot the largest $\tilde{k}_{y}\tilde{k}_{z}$ cross sections ($\tilde{k}_{x}=0$) of bulk Fermi surfaces, which are identical for the two sides of the junction. (d) Josephson energy phase relations for states at the representative transverse momenta (in color) and states at the same values of $\tilde{k}_{z}$ but $k_{y}=\pm 0.1,\pm 0.2$ (in gray). Here, the system size and other parameters are the same as those taken in Fig. 3.
@@
We study a tight-binding model of the Josephson junction between two identical $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductors on a cubic lattice, similar to that discussed in Sec. IV.1.2, but with the junction now aligned along the $x$ direction, as shown schematically in Fig. 5 (a). Due to the mirror $M_{z}$ symmetry of the system, there are identical contributions to the Josephson energy phase relation from $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$. Thus, we first examine states at five representative transverse momentum $\tilde{\mathbf{k}}_{i}=(0,\tilde{k}_{i})$ ($i=1,\dots,5$) within $\mathrm{FS}_{1}$ in the $k_{y}k_{z}$ plane, as shown in Fig. 5 (c). At each conserved $\tilde{\mathbf{k}}_{i}$, the effective 1D Josephson junction along $x$ direction hosts midgap states localized at the junction interface, which contribute to $4\pi$-periodic Josephson energy phase relation, as shown in Fig. 5 (d). As the magnitude of $\tilde{k}_{z}$ increases towards the locations of pairing gap nodes, the effective pairing amplitude, $\tilde{\Delta}(\tilde{k}_{z})$, decreases. Furthermore, at phase difference $\phi=\pi$, the junction interface plays the role of a domain wall, where the pairing order changes its sign at the junction barrier. For each representative transverse momentum at $\tilde{k}_{y}=0$, the pairing order is odd under mirror $M_{x}$ reflection, ${\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}(\tilde{k}_{x},\tilde{k}_{y}=0,\tilde{k}_{z})}={-\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}(-\tilde{k}_{x},\tilde{k}_{y}=0,\tilde{k}_{z})}$. Together with the BdG particle-hole symmetry, this system is analogous to the Jackiw-Rebbi model [40] and exhibits zero energy Andreev states localized at junction interface when $\phi=\pi$. When fermion parity is conserved at $\phi=\pi$, the total energy phase relation is $4\pi$-periodic and leads to Josephson current phase relation of the same periodicity, shown in Fig. 6 (a) in red. As a result, when a nonzero external magnetic flux is varied through the junction interface, the Josephson Fraunhofer pattern exhibits lifted nodes at odd integer flux quanta, shown in red in Fig. 6 (b). The conserved fermion parity can be achieved experimentally, for example, by taking a dynamic measurement of the Josephson current while maintaining a fixed voltage to prevent parity transitions [41, 42, 43].
@@
The existence of $4\pi$-periodic Josephson energy phase relation in this single Josephson junction oriented along the $x$ direction stems from the local spatial symmetry property of the monopole pairing order, which contains nonzero component of $p_{x}$-wave pairing, or say, nonzero components of $l_{z,\mathrm{loc}}=\pm 1$. Furthermore, due to spin-orbit coupling from Weyl band Hamiltonian, we examine the pairing order under combined spin and spatial rotations. The $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor is invariant under simultaneous spin and spatial rotations; hence, the local $z$-component angular momentum, $l_{z,\mathrm{loc}}$, in momentum space and of spin, $s_{z}$, of the pairing order satisfies $l_{z,\mathrm{loc}}+s_{z}=0$. However, the existence of $4\pi$-periodic Josephson current constrains $l_{z,\mathrm{loc}}\neq 0$, i.e. the monopole superconducting order has the symmetry of a chiral superconducting order locally in momentum space. Specifically, at conserved momentum $\tilde{k}_{z}$, the above junction can only be described in analogy to that between two superconductors with spin-polarized chiral $p$-wave pairing order, in which the Cooper pair is locally described by $(p_{x}+ip_{y})|\downarrow\downarrow\rangle$ or $(p_{x}-ip_{y})|\uparrow\uparrow\rangle$ pairing, satisfying $l_{z,\mathrm{loc}}=-s_{z}$. Later, in Sec. VI, we show that the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconducting order is indeed unique and different from these chiral superconductors globally. Nonetheless, though the monopole superconducting order is inherently different from any pairing based on spherical harmonic symmetry, at given $\tilde{k}_{z}$ one can describe the symmetry of the monopole pairing order locally in analogy to that of a 2D chiral superconducting pairing order. As such, for given conserved $\tilde{k}_{z}$, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ superconductor can be described locally in momentum space as having chiral $p$-wave symmetry.
@@
If the symmetry of the spin-orbit coupled system is not known a priori, we additionally examine a second Josephson junction to demonstrate the local symmetry of pairing order. Consider the junction geometry in which the monopole superconductor to the left of the junction barrier is rotated by an angle $\theta$ about the $z$ axis, as shown in Fig. 5 (b). If the Josephson energy phase relation is invariant under the rotation, exhibiting $4\pi$-periodicity with no phase shift, the pairing order lies in the zero total angular momentum sector with a nonzero orbital angular momentum $l_{z,\mathrm{loc}}$. We check this numerically using the tight-binding model, where the four-fold lattice rotation takes discrete values of $\theta=n\pi/2$ with $n\in\mathbb{Z}$, and obtain the same Josephson energy phase relation as in Fig. 5 (d). Then, following the same reasoning as in the paragraph above, this additional rotated junction demonstrates the local chiral $p$-wave symmetry of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole pairing order.

### v2

@@ IV.3 Generalization to monopole pairing orders with higher monopole charges
We further demonstrate the probe of global angular momentum for monopole superconductors in different topological sectors characterized by higher monopole charges, $|q_{p}|>1$. Although the topological sector is different, the same symmetry principles to extract $l_{z,\mathrm{gob}}$ outlined in Sec. III.2 still hold. We begin with the general model of the monopole superconductor in Eq. (13) but now consider a multi-Weyl semimetal with inter-Fermi surface pairing between two parity-related Fermi surfaces. For three-dimensional Weyl semimetals on a lattice with $n$-fold rotational symmetry ($n=2,3,4,6$), there can exist single, double, or triple Weyl nodes Fang et al. (2012); Tsirkin et al. (2017). The lattice point group symmetry sets an upper bound to the chirality of the Weyl node, with $|\nu|\leq[n/2]$, and similarly for the pair monopole charge, $|q_{p}|\leq n/2$.
@@
where $\sigma_{\pm}=\left(\sigma_{x}\pm i\sigma_{y}\right)/2$, and is related to the other Fermi surface by parity, $\mathcal{H}_{\mathrm{kin},1}^{(\nu=-3)}(\mathbf{k})=\sigma_{z}\mathcal{H}_{\mathrm{kin},2}^{(\nu=+3)}(-\mathbf{k})\sigma_{z}$. The helical eigenstate at $\mathrm{FS}_{1}$ is $\chi_{1,+}(\tilde{\mathbf{k}})=(\cos({\theta_{h_{1}}}/{2})e^{-i(\lambda-1)\varphi_{h_{1}}/2},\sin({\theta_{h_{1}}}/{2})e^{-i(\lambda+1)\varphi_{h_{1}}/2})$, and at $\mathrm{FS}_{2}$, it is $\chi_{2,+}(\tilde{\mathbf{k}})=\sigma_{z}\chi_{1,+}(-\tilde{\mathbf{k}})$. Here, consistent with previous notation, $\lambda=\pm 1$ denotes the choice of gauge, and $\theta_{h_{1}}$ and $\varphi_{h_{1}}$ are the polar and azimuthal angles defined in $\mathbf{h}_{1}^{(\nu=-3)}$ spherical coordinates by $\mathbf{h}_{1}^{(\nu=-3)}\equiv|h_{1}^{(\nu=-3)}|(\sin\theta_{h_{1}}\cos\varphi_{h_{1}},\sin\theta_{h_{1}}\sin\varphi_{h_{1}},\cos\theta_{h_{1}})$. They are related to momentum space polar and azimuthal angles by $\sin\theta_{h_{1}}=\sin^{3}\theta_{\tilde{k}}/{(\cos^{2}\theta_{\tilde{k}}+\sin^{6}\theta_{\tilde{k}})^{1/2}}$ and $\varphi_{h_{1}}=-3\varphi_{\tilde{k}}$. We study the case of inter-Fermi surface $s$-wave pairing in Eq. (13), where $\Delta_{\mathrm{inter}}(\mathbf{k})=i\Delta_{0}\sigma_{y}$, with $\Delta_{0}$ being the pairing amplitude.
In the weak-coupling regime, we obtain the low-energy band-projected pairing order from Eq. (16) which takes the following form:
@@
We now examine the Josephson coupling between the above $q_{p}=3$ monopole superconductor and a spin singlet superconductor, following the methods introduced in Sec. III.1. For comparison, we consider the case where there is bare spin-independent tunneling at the junction barrier. The form factor for this system evaluates to
@@
where $\bm{\Omega}_{h_{1}}=(\theta_{h_{1}},\varphi_{h_{1}})$ is the solid angle in $\mathrm{h}_{1}^{(\nu=3)}$-space and $\Delta_{\mathrm{MSC}}^{(q_{p}=3}(\tilde{\mathbf{k}})$ is the effective pairing order of the $q_{p}=3$ monopole superconductor. Similar to the case of the $q_{p}=-1$ monopole superconductor, whose form factor is given in Eq. (27), the half-integer monopole harmonics arise from the effective spin-dependent tunneling arising from the spin-orbit coupling in the triple Weyl semimetal. The above form factor is non-singular and can be used to determine the first-order Josephson coupling of this system.
@@
where $\Delta_{L,0}$ and $\Delta_{R,0}$ are the pairing amplitudes of the $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ monopole superconductor respectively. $\mathfrak{F}_{s-\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ itself is invariant under rotation $R_{z}$ and parity, as both the $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ monopole superconductor both form irreps of $l_{z}=0$. As discussed in Sec. IV.1.1, this leads to nonvanishing first-order Josephson current. Hence, following the junction design principles introduced in Sec. III.2, we may similarly design probes of monopole pairing order in different topological sectors.
@@ V Josephson junctions to probe $l_{z,\mathrm{loc}}$ of the monopole superconducting order
We now explore how to design Josephson junctions to probe local angular momentum $l_{z,\mathrm{loc}}$ of monopole superconducting order, following the second junction design introduced in Sec. III.2. We consider two identical monopole superconductors forming a Josephson junction oriented along, for example, the $x$ direction, which is perpendicular to the high-symmetry rotational axis along $z$. As discussed in Sec. III.2, the total Josephson current can be regarded as a superposition of contributions from a family of Josephson junctions between two $2$D chiral superconductors, corresponding to different values of conserved momentum $k_{z}$. It has been studied extensively in, for example, Refs. 23; 29 that a planar junction between two chiral superconductors can host $4\pi$-periodic Josephson current arising from the tunneling of chiral boundary modes. Therefore, by studying the Josephson energy phase relation in our system, we can obtain the information about local symmetry of monopole pairing order. In the continuum model of a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, its pairing order, described by Eq. (18), exhibits local chiral $p$-wave pairing with $l_{z,\mathrm{loc}}=\mp 1$ for a given value of $\tilde{k}_{z}$,
@@
Figure 5: (a) Single Josephson junction along the $x$ direction between two identical monopole superconductors with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ and the same crystal orientation. (b) Same Josephson junction as in (a), only now that the superconductor at the left (yellow) is rotated by an angle $\theta$ about the $z$ axis pointing out of plane. (c) Bulk Fermi surface $\mathrm{FS}_{1}$ of the doped Weyl semimetal on a cubic lattice. Representative momenta $\tilde{\mathbf{k}}_{i=1,\cdots,5}$ are shown, with $\tilde{k}_{y}=0$ and $\tilde{k}_{z}=-0.2,-0.1,0.0,0.1,0.15$, respectively. As a reference, we plot the largest $\tilde{k}_{y}\tilde{k}_{z}$ cross sections ($\tilde{k}_{x}=0$) of bulk Fermi surfaces, which are identical for the two sides of the junction. (d) Josephson energy phase relations for states at the representative transverse momenta (in color) and states at the same values of $\tilde{k}_{z}$ but $k_{y}=\pm 0.1,\pm 0.2$ (in gray). Here, the system size and other parameters are the same as those taken in Fig. 3.
@@
We study a tight-binding model of the Josephson junction between two identical $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductors on a cubic lattice, similar to that discussed in Sec. IV.1.2, but with the junction now aligned along the $x$ direction, as shown schematically in Fig. 5(a). Due to the mirror $M_{z}$ symmetry of the system, there are identical contributions to the Josephson energy phase relation from $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$. Thus, we first examine states at five representative transverse momentum $\tilde{\mathbf{k}}_{i}=(0,\tilde{k}_{i})$ ($i=1,\dots,5$) within $\mathrm{FS}_{1}$ in the $k_{y}k_{z}$ plane, as shown in Fig. 5(c). At each conserved $\tilde{\mathbf{k}}_{i}$, the effective $1$D Josephson junction along $x$ direction hosts midgap states localized at the junction interface, which contribute to $4\pi$-periodic Josephson energy phase relation, as shown in Fig. 5(d). As the magnitude of $\tilde{k}_{z}$ increases towards the locations of pairing gap nodes, the effective pairing amplitude, $\tilde{\Delta}(\tilde{k}_{z})$, decreases. Furthermore, at phase difference $\phi=\pi$, the junction interface plays the role of a domain wall, where the pairing order changes its sign at the junction barrier. For each representative transverse momentum at $\tilde{k}_{y}=0$, the pairing order is odd under mirror $M_{x}$ reflection, ${\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}(\tilde{k}_{x},\tilde{k}_{y}=0,\tilde{k}_{z})}={-\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}(-\tilde{k}_{x},\tilde{k}_{y}=0,\tilde{k}_{z})}$. Together with the BdG particle-hole symmetry, this system is analogous to the Jackiw-Rebbi model Jackiw and Rebbi (1976) and exhibits zero energy Andreev states localized at junction interface when $\phi=\pi$. When fermion parity is conserved at $\phi=\pi$, the total energy phase relation is $4\pi$-periodic and leads to Josephson current phase relation of the same periodicity, shown in Fig. 6(a) in red. As a result, when a nonzero external magnetic flux is varied through the junction interface, the Josephson Fraunhofer pattern exhibits lifted nodes at odd integer flux quanta, shown in red in Fig. 6(b). The conserved fermion parity can be achieved experimentally, for example, by taking a dynamic measurement of the Josephson current while maintaining a fixed voltage to prevent parity transitions Kurter et al. (2015); Abboud et al. (2022); Yue et al. (2024).
@@
The existence of $4\pi$-periodic Josephson energy phase relation in this single Josephson junction oriented along the $x$ direction stems from the local spatial symmetry property of the monopole pairing order, which contains nonzero component of $p_{x}$-wave pairing, or say, nonzero components of $l_{z,\mathrm{loc}}=\pm 1$. Furthermore, due to spin-orbit coupling from Weyl band Hamiltonian, we examine the pairing order under combined spin and spatial rotations. The $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor is invariant under simultaneous spin and spatial rotations; hence, the local $z$-component angular momentum, $l_{z,\mathrm{loc}}$, in momentum space and of spin, $s_{z}$, of the pairing order satisfies $l_{z,\mathrm{loc}}+s_{z}=0$. However, the existence of $4\pi$-periodic Josephson current constrains $l_{z,\mathrm{loc}}\neq 0$, i.e. the monopole superconducting order has the symmetry of a chiral superconducting order locally in momentum space. Specifically, at conserved momentum $\tilde{k}_{z}$, the above junction can only be described in analogy to that between two superconductors with spin-polarized chiral $p$-wave pairing order, where the Cooper pair is locally described by $(p_{x}+ip_{y})|\downarrow\downarrow\rangle$ or $(p_{x}-ip_{y})|\uparrow\uparrow\rangle$ pairing, satisfying $l_{z,\mathrm{loc}}=-s_{z}$. Later, in Sec. VII, we show that the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconducting order is indeed unique and different from these chiral superconductors globally. Nonetheless, though the monopole superconducting order is inherently different from any pairing based on spherical harmonic symmetry, at given $\tilde{k}_{z}$, one can describe the symmetry of the monopole pairing order locally in analogy to that of a $2$D chiral superconducting pairing order. As such, for given conserved $\tilde{k}_{z}$, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ superconductor can be described locally in momentum space as having chiral $p$-wave symmetry. Furthermore, the qualitative features of the Josephson current phase relation in Fig. 6(a), specifically the periodicity, are unchanged in the long junction limit, as discussed in Sec. VI.
@@
If the symmetry of the spin-orbit coupled system is not known a priori, we additionally examine a second Josephson junction to demonstrate the local symmetry of pairing order. Consider the junction geometry where the monopole superconductor to the left of the junction barrier is rotated by an angle $\theta$ about the $z$ axis, as shown in Fig. 5(b). If the Josephson energy phase relation is invariant under the rotation, exhibiting $4\pi$-periodicity with no phase shift, the pairing order lies in the zero total angular momentum sector with a nonzero orbital angular momentum $l_{z,\mathrm{loc}}$. We check this numerically using the tight-binding model, where the fourfold lattice rotation takes discrete values of $\theta=n\pi/2$ with $n\in\mathbb{Z}$, and obtain the same Josephson energy phase relation as in Fig. 5(d). Then, following the same reasoning as in the paragraph above, this additional rotated junction demonstrates the local chiral $p$-wave symmetry of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole pairing order.
@@
VI Long Josephson junctions
Figure 7: (a) SNS Josephson junction along the $z$ direction between and $s$-wave and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, analogous to that of Fig. 2(a) but consisting of a normal metal (NM) region of length $L$ shown in gray. (b) Josephson critical current as a function of $L$ in (a), normalized to the critical current for the junction of $L=1$. Parameters are identical to those in Fig. 3. (c) SNS Josephson junction along the $x$ direction between identical $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductors, analogous to that of Fig. 5(a) but consisting of a normal metal (NM) region of length $L$ shown in gray. (d) Josephson critical current as a function of $L$ in (c), normalized to the critical current for the junction of $L=1$. Parameters are identical to those in Fig. 5.
In this section, we explore the Josephson effect in long superconductor-normal metal-superconductor (SNS) junctions using tight-binding models, where the width of the normal metal region exceeds the superconducting coherence length. Our numerical calculations show that both the nonzero Josephson coupling and the $2\pi$- and $4\pi$-periodic Josephson current phase relations, as discussed in Secs. IV and V, are not exclusive to the short junction regime. Instead, these fundamental features arise from the symmetry of the superconducting order and the presence of topological boundary states, which remain robust for varying junction lengths as long as ballistic transport is maintained in the normal metal regime. However, the critical Josephson current magnitude decreases with increasing junction length.
We consider the tight-binding model Hamiltonian, analogous to the setup in Appendix D, $H_{\mathrm{JJ}}(\mathbf{k}_{\parallel})=H_{\mathrm{BdG},L}(\mathbf{k}_{\parallel})+H_{\mathrm{link}}(\mathbf{k}_{\parallel})+H_{\mathrm{BdG},R}(\mathbf{k}_{\parallel}).$ Here, the junction is oriented along the $\hat{\mathbf{n}}$ direction ($\hat{\mathbf{n}}=\hat{x},\hat{z}$), and transverse momentum $\mathbf{k}_{\parallel}$ is conserved. The bulk superconductor BdG Hamiltonians on the left and right sides follow the form of Eq. (82). For the normal metal region, the tight-binding Hamiltonian is defined as
$H_{\mathrm{link}}(\mathbf{k}_{\parallel})=\sum_{\langle n,n^{\prime}\rangle}\sum_{\sigma=\uparrow,\downarrow}t_{0}\delta_{n,n^{\prime}+1}c^{\dagger}_{n,\mathbf{k}_{\parallel};\sigma}c_{n^{\prime},\mathbf{k}_{\parallel};\sigma}+\mathrm{h.c.},$ (41)
where the nearest neighbor sum $\langle n,n^{\prime}\rangle$ runs from site $N$ to site $N+L$ along $\hat{\mathbf{n}}$ at the two ends of the normal metal region. For $L=1$, this reduces to Eq. (33). Here, we assume spin-independent and transverse momentum-independent nearest-neighbor hopping with constant amplitude $t_{0}=-0.5$.
We first study a long Josephson junction oriented along the $z$ direction, between an $s$-wave and a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, as shown schematically in Fig. 7(a). The Fermi velocities for the left and right superconductors are $v_{F,L}=1.00$ and $v_{F,R}=0.90$, with a pairing amplitude $\Delta_{0}=0.02$ which corresponds to coherence lengths $\xi_{L}\approx 50$ and $\xi_{R}\approx 45$ in the unit of lattice constant. As the length of normal metal region $L$ increases from $1$ to $151$, the first-order Josephson coupling remains nonvanishing, with a $2\pi$ periodic Josephson energy phase relation similar to that shown in Fig. 3(c) for the short junction studied in Sec. IV.1.2. The critical Josephson current $I_{C}$, shown in Fig. 7(b), initially increases for $L=11$ and then decays exponentially as $L$ increases from $11$ to $151$.
Next, we examine an SNS junction along the $x$ direction between identical $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductors, as described in Sec. V, with Fermi velocity $v_{F}=0.51$ and coherence length $\xi\approx 25$. We find nonvanishing Josephson current and $4\pi$-periodic Josephson energy phase relation, which is qualitatively the same as that shown in Fig. 6(a), as the junction length increases from $L=1$ to $L=151$. The zero-energy Majorana bound states still contribute $4\pi$ periodic Josephson current, as discussed in Sec. V. The Josephson critical current decays exponentially for increasing $L$, as presented in Fig. 7(d).
Overall, for both junctions in Figs. 7(a) and 7(b), the qualitative properties of Josephson current and its current phase relation remain unchanged despite an exponential decay in the critical current for long junctions. These results affirm that even in the long SNS junction geometry, our design principles remain valid to identify the monopole superconducting order via probing the global and local angular momenta of the monopole superconducting order through the Josephson coupling.